# Generalized Spherical Neural Operators: Green’s Function Formulation

### v1

@@ H V D Average H V D Average
ClimaX (nguyen2023climax) $5.27\pm 0.16$ $222\pm 9.00$ $1000\pm 0.13$ $409\pm 3.00$ $5.87\pm 0.19$ $324\pm 11.7$ $1000\pm 0.18$ $443\pm 3.90$
FourCastNet (pathak2022fourcastnet) $3.40\pm 0.12$ $187\pm 4.86$ $716\pm 7.69$ $302\pm 2.99$ $3.56\pm 0.08$ $319\pm 6.15$ $737\pm 8.95$ $353\pm 3.45$
SFNONet (bonev2023spherical) $1.39\pm 0.07$ $145\pm 11.1$ $295\pm 8.87$ $147\pm 4.89$ $1.68\pm 0.12$ $229\pm 15.6$ $353\pm 13.5$ $195\pm 6.86$
GSHNet (Ours) $\mathbf{1.26\pm 0.07}$ $\mathbf{134\pm 9.30}$ $\mathbf{273\pm 5.91}$ $\mathbf{136\pm 3.73}$ $\mathbf{1.49\pm 0.09}$ $\mathbf{201\pm 11.9}$ $\mathbf{326\pm 12.1}$ $\mathbf{176\pm 5.71}$
@@ 5 Experiments
We compare GSNO and GSHNet to other state-of-the-art methods under identical experimental setup and configuration (liu2024evaluation), e.g., Transformer-based ClimaX (nguyen2023climax), FNO-based FourCastNet (pathak2022fourcastnet), and SFNO-based SFNONet (bonev2023spherical) (detailed in Appendix C.1). The models are compared on diffusion magnetic resonance imaging (dMRI) modeling and two autoregressive spherical datasets. We also conduct ablation experiments to further verify the effectiveness of the proposed GSNO block and GSHNet structure. All experiments are implemented through PyTorch on 16GB A5000 GPUs.
@@ 5.1 Spherical Shallow Water Equations
Spherical Shallow Water Equations (SSWE) form a nonlinear hyperbolic PDEs system that model the motion of thin-layer fluids on a rotating sphere. The core underlying assumption is the shallow water approximation, where the vertical scale of the fluid layer is much smaller than the horizontal scale (bonev2018discontinuous). Following (bonev2018discontinuous; bonev2023spherical), we generate the SSWE simulation using a classical spectral solver (giraldo2001spectral). This dataset is well-suited to verify our model due to its spherical geometric characteristics (Appendix C.2, spatial resolution of $256\times 512$, time step of 60s, 3 channel dimensions: geopotential height (H), vorticity (V), and divergence (D)). All comparisons use the identical dataset and setting: 50 epochs containing 256 samples each, batch size of 16, Adam optimizer with learning rate of $1\times 10^{-3}$, and spherical weighted mean relative loss (details in Appendix C.1). Models are tested on the dataset generated from 50 initial conditions (50 samples) and evaluated using MRE for each variable. The results (Table 1) show that our method achieves the best performance on all variables and time scales, where the average increase is 7.5% at 5h and 9.7% at 10h. Besides, we conduct additional verifications at different resolution (128 × 256) and the results are summarized in the Table 6 in Appendix. For predicting geopotential height (Figure 3), SFNONet and FourCastNet show more errors in the zoomed-in region, compared to our model. ClmaX is not presented due to its poor results.
@@
Figure 4: Predicted V600 samples of different methods. Each image set shows the main prediction (left), a zoomed-in region (top-right), and the residual (Res) to the ground truth (bottom-right, darker is better). The performance is evaluated using ACC(×10-2)/MSE (×10-2).
5.2 Weather Forecasting
To further assess the performance of our methods, we utilize WeatherBench (rasp2020weatherbench), a widely used autoregressive global weather prediction benchmark that offers data at multiple spatial resolutions. We select the dataset with a spatial resolution of $5.625^{\circ}$ ( 32 × 64) and a temporal resolution of 1 hour for evaluation. The training period spans 1979-2015, validation is conducted on data from 2016, and testing covers 2017-2018. The full dataset includes 24 meteorological variables, and we select six key variables as prediction targets (liu2024evaluation) (details in Appendix C.3). All models are under the same settings: batch size 128, 50 epochs, loss function of MRE, Adam optimizer with learning rate of $4\times 10^{-4}$. To evaluate multi-scale forecasting ability, we perform prediction at 24, 72, 120 autoregressive steps. Performance is measured using anomaly correlation coefficient (ACC) and latitude-weighted mean square error (MSE) (nguyen2023climax) (details in Appendix C.3), reported in Table 2 and 8 (details in Appendix C.3). Besides, we conduct additional verifications at different resolution (64 × 128) and the results are summarized in the Table 6 in Appendix. Our method achieves the best results on all variables and time scales, especially in the predictions for the third and fifth days. The wind velocity predictions (V600) and their residual plots (darker plots are better) also demonstrate the accuracy and stability of our method in Figure 4.
To further evaluate superiority of the spectral kernel $G_{\theta_{2}}^{2}(l,m)$ in GSNO, we construct extra operator comparison experiments, with all other architectures and settings identical for fair comparison. The results in the Table LABEL:PE_vs clearly demonstrate that our GSNO consistently and significantly outperforms the SFNOs with higher capacity and spatial position-encoding SNO ($G_{\theta}(x,y)$) across all forecast lead times (details in Appendix C.3).
5.3 Diffusion MRI modelling of brain microstructure
Modeling brain microstructure with diffusion MRI is challenging due to sparse, anisotropic measurements across spherical shells (van2013wu; jeurissen2014multi; zeng2022fod); We choose dMRI-based Fiber Orientation Distribution (FOD) angular super-resolution (zeng2022fod; snoussi2025equivariant) as the specific task for evaluating model performance on anisotropic system; see Appendix C.4 for challenge, motivation and preprocessing details. To further verify GSNO, we adopt the same architecture as FODNet (zeng2022fod), only changing the convolutional layers, adapting it to SFNOs to obtain FOD-SFNO, and GSNOs to obtain FOD-GSNO. ACC is used for evaluation, which stands for Angular Correlation Coefficient, a standard metric for measuring the angular similarity between two spherical functions (zeng2022fod). The results in Table 4 show that GSNO consistently outperforms other models, demonstrating strong generalization under irregular sampling and data sparsity. Figure 5 shows the region of interest taken from the cingulate gyrus. Compared with the obviously false positive predictions by other methods, the proposed GSNO achieves better prediction of fiber orientation and density.
@@ 1 Day Forecast
Climax (nguyen2023climax) - 5.4 M 96.2 ± 2.01 91.7 ± 0.67 91.4 ± 0.75 92.5 ± 0.67 90.4 ± 0.82 96.2 ± 1.38 93.1 ± 0.47
FourCastNet (pathak2022fourcastnet) FNO 5.3 M 96.0 ± 2.12 90.5 ± 0.75 90.0 ± 0.85 91.9 ± 0.78 90.1 ± 0.95 95.5 ± 1.62 92.3 ± 0.52
SFNONet (bonev2023spherical) SFNO 5.3 M 95.2 ± 2.53 81.7 ± 1.81 82.1 ± 1.78 86.4 ± 1.50 84.0 ± 2.02 94.1 ± 2.39 87.3 ± 0.83
GSHNet (Ours) GSNO 4.0 M 96.8 ± 2.00 88.3 ± 1.12 88.5 ± 1.10 90.1 ± 1.07 89.3 ± 1.29 96.0 ± 1.78 91.5 ± 0.59
SFNONet (bonev2023spherical) GSNO 5.5 M 95.6 ± 2.37 84.3 ± 1.62 84.1 ± 1.70 87.6 ± 1.35 85.8 ± 1.66 94.7 ± 2.21 88.7 ± 0.76
GSHNet SFNO 3.7 M 96.0 ± 2.21 86.9 ± 1.47 85.8 ± 1.50 88.3 ± 1.29 86.9 ± 1.42 95.3 ± 1.92 89.9 ± 0.68
@@ 3 Day Forecast
Climax (nguyen2023climax) - 5.4 M 90.0 ± 5.69 58.1 ± 5.08 56.3 ± 5.12 66.8 ± 4.62 57.5 ± 5.30 83.4 ± 8.04 68.7 ± 2.35
FourCastNet (pathak2022fourcastnet) FNO 5.3 M 88.7 ± 6.59 56.3 ± 4.83 54.4 ± 4.72 66.0 ± 4.40 55.2 ± 5.40 81.8 ± 8.49 67.1 ± 2.41
SFNONet (bonev2023spherical) SFNO 5.3 M 90.9 ± 5.32 58.0 ± 5.39 55.2 ± 5.28 67.2 ± 4.59 57.8 ± 5.66 83.4 ± 8.30 68.8 ± 2.40
GSHNet (Ours) GSNO 4.0 M 92.5 ± 4.63 61.7 ± 4.68 60.8 ± 4.72 70.2 ± 4.21 63.6 ± 5.01 85.8 ± 7.46 72.4 ± 2.13
SFNONet (bonev2023spherical) GSNO 5.5 M 91.4 ± 4.97 59.8 ± 5.03 58.7 ± 4.95 68.4 ± 4.42 60.1 ± 5.40 84.3 ± 7.92 70.5 ± 2.27
GSHNet SFNO 3.7 M 91.8 ± 4.82 60.6 ± 4.90 57.2 ± 5.06 68.7 ± 4.44 61.0 ± 5.22 84.6 ± 7.85 70.7 ± 2.24
@@ 5 Day Forecast
Climax (nguyen2023climax) - 5.4 M 84.0 ± 9.63 28.2 ± 9.35 21.3 ± 8.23 35.7 ± 10.0 14.2 ± 8.57 66.3 ± 18.4 41.6 ± 4.59
FourCastNet (pathak2022fourcastnet) FNO 5.3 M 83.0 ± 11.3 29.2 ± 8.68 22.5 ± 7.50 36.2 ± 10.5 13.8 ± 6.47 64.5 ± 19.9 41.5 ± 4.74
SFNONet (bonev2023spherical) SFNO 5.3 M 86.9 ± 7.76 35.8 ± 8.74 29.5 ± 9.28 45.3 ± 9.33 26.5 ± 9.47 71.2 ± 16.0 49.2 ± 4.27
GSHNet (Ours) GSNO 4.0 M 88.2 ± 7.21 40.3 ± 8.32 35.2 ± 8.37 47.7 ± 8.71 30.7 ± 9.29 73.0 ± 14.8 52.5 ± 3.99
SFNONet (bonev2023spherical) GSNO 5.5 M 87.7 ± 7.68 38.6 ± 8.46 33.5 ± 8.72 47.1 ± 8.92 28.9 ± 9.35 72.1 ± 15.3 51.3 ± 4.11
GSHNet SFNO 3.7 M 87.9 ± 7.40 38.2 ± 8.62 32.3 ± 8.85 46.3 ± 9.11 28.5 ± 9.36 72.4 ± 15.1 50.9 ± 4.10
Table 3: Performance comparison of different spherical operator models. Bold is the best result.
@@
Methods
SSMT-CSD
FOD-Net
FOD-SFNO
ESCNN
FOD-GSNO
@@ 5.4 Ablation Experiments and computational costs
To further evaluate the contributions of GSNO and GSHNet , we conduct ablation studies to replace the GSNO block with a classic SFNO block in the GSHNet or replace the classic SFNO block with GSNO in the SFNONet, leading to performance degradation across all variables and time horizons (Table 2 and 8 (Appendix C.3)). This confirms the effectiveness of GSNO and the multi-scale architecture of GSHNet. Further, we observe more pronounced performance gains from GSNO blocks in longer-term predictions, suggesting that GSNO plays an increasingly important role in modelling long-range dynamics. To demonstrate the efficiency comparison, we provide the parameters of different models in Table 2. Detailed computational costs are in Appendix C.6.
@@ 6 Conclusion
We have introduced a rigorous theoretical framework for spherical neural operators, grounded in the formulation of designable Green’s function on the sphere, and extended to the spherical harmonic domain. This framework establishes the foundation for our proposed Green’s Function Formulation, which enabled the principled incorporation of complex constraints on spherical operator learning. Building on this framework, we design an absolute and relative position-dependent Green’s function, yielding a novel spherical learning method that captures complex physical constraints for asymmetric conditions and localized phenomena. The proposed formulation leads to GSNO, designed to handle a broader range of real-world scenarios. Further, we design a multi-scale architecture, GSHNet, incorporating geometrically adaptive up-and-down sampling to enhance interactions across resolutions on the sphere. Both theoretical derivation and extensive experiments consistently demonstrate the superiority of our approaches over other state-of-the-art methods.

### v2

@@ H V D Average H V D Average
ClimaX (Nguyen et al., 2023) $5.27\pm 0.16$ $222\pm 9.00$ $1000\pm 0.13$ $409\pm 3.00$ $5.87\pm 0.19$ $324\pm 11.7$ $1000\pm 0.18$ $443\pm 3.90$
FourCastNet (Pathak et al., 2022) $3.40\pm 0.12$ $187\pm 4.86$ $716\pm 7.69$ $302\pm 2.99$ $3.56\pm 0.08$ $319\pm 6.15$ $737\pm 8.95$ $353\pm 3.45$
SFNONet (Bonev et al., 2023) $1.39\pm 0.07$ $145\pm 11.1$ $295\pm 8.87$ $147\pm 4.89$ $1.68\pm 0.12$ $229\pm 15.6$ $353\pm 13.5$ $195\pm 6.86$
SHNet (Ours) $\mathbf{1.26\pm 0.07}$ $\mathbf{134\pm 9.30}$ $\mathbf{273\pm 5.91}$ $\mathbf{136\pm 3.73}$ $\mathbf{1.49\pm 0.09}$ $\mathbf{201\pm 11.9}$ $\mathbf{326\pm 12.1}$ $\mathbf{176\pm 5.71}$
@@ 5 Experiments
We compare GSNO and SHNet to other state-of-the-art methods under identical experimental setup and configuration (Liu et al., 2024), e.g., Transformer-based ClimaX (Nguyen et al., 2023), FNO-based FourCastNet (Pathak et al., 2022), and SFNO-based SFNONet (Bonev et al., 2023) (detailed in Appendix C.1). The models are compared on diffusion magnetic resonance imaging (dMRI) modeling and two autoregressive spherical datasets. We also conduct ablation experiments to further verify the effectiveness of the proposed GSNO block and SHNet structure. All experiments are implemented through PyTorch on 16GB A5000 GPUs.
@@ 5.1 Spherical Shallow Water Equations
Spherical Shallow Water Equations (SSWE) form a nonlinear hyperbolic PDEs system that model the motion of thin-layer fluids on a rotating sphere. The core underlying assumption is the shallow water approximation, where the vertical scale of the fluid layer is much smaller than the horizontal scale (Bonev et al., 2018). Following (Bonev et al., 2018; 2023), we generate the SSWE simulation using a classical spectral solver (Giraldo, 2001). This dataset is well-suited to verify our model due to its spherical geometric characteristics (Appendix C.2, spatial resolution of $256\times 512$, time step of 60s, 3 channel dimensions: geopotential height (H), vorticity (V), and divergence (D)). All comparisons use the identical dataset and setting: 50 epochs containing 256 samples each, batch size of 16, Adam optimizer with learning rate of $1\times 10^{-3}$, and spherical weighted mean relative loss (details in Appendix C.1). Models are tested on the dataset generated from 50 initial conditions (50 samples) and evaluated using MRE for each variable. The results (Table 1) show that our method achieves the best performance on all variables and time scales, where the average increase is 7.5% at 5h and 9.7% at 10h. Besides, Besides, we conduct additional verifications under the exact same setting as SFNONet and at different resolution ($128\times 256$) for reference, the results are summarized in the Table 6 and LABEL:same_as_SFNO in Appendix. For the predicted geopotential height (Figure 3), SFNONet and FourCastNet show more errors in the zoomed-in region, compared to our model. ClmaX is not presented due to its poor results.
Figure 3: Predicted geopotential height (H) at 5h and 10h from different methods. Zoom in and provide the absolute residual (Res) to the ground truth for clarity (darker is better).
@@
Figure 4: Predicted V600 samples of different methods. Each image set shows the main prediction (left), a zoomed-in region (top-right), and the absolute residual (Res) to the ground truth (bottom-right, darker is better). The performance is evaluated using ACC(×10-2)/MSE (×10-2).
@@ 1 Day Forecast
Climax (Nguyen et al., 2023) - 5.4 M 96.2 ± 2.01 91.7 ± 0.67 91.4 ± 0.75 92.5 ± 0.67 90.4 ± 0.82 96.2 ± 1.38 93.1 ± 0.47
FourCastNet (Pathak et al., 2022) FNO 5.3 M 96.0 ± 2.12 90.5 ± 0.75 90.0 ± 0.85 91.9 ± 0.78 90.1 ± 0.95 95.5 ± 1.62 92.3 ± 0.52
SFNONet (Bonev et al., 2023) SFNO 5.3 M 95.2 ± 2.53 81.7 ± 1.81 82.1 ± 1.78 86.4 ± 1.50 84.0 ± 2.02 94.1 ± 2.39 87.3 ± 0.83
SHNet (Ours) GSNO 4.0 M 96.8 ± 2.00 88.3 ± 1.12 88.5 ± 1.10 90.1 ± 1.07 89.3 ± 1.29 96.0 ± 1.78 91.5 ± 0.59
SFNONet (Bonev et al., 2023) GSNO 5.5 M 95.6 ± 2.37 84.3 ± 1.62 84.1 ± 1.70 87.6 ± 1.35 85.8 ± 1.66 94.7 ± 2.21 88.7 ± 0.76
SHNet SFNO 3.7 M 96.0 ± 2.21 86.9 ± 1.47 85.8 ± 1.50 88.3 ± 1.29 86.9 ± 1.42 95.3 ± 1.92 89.9 ± 0.68
@@ 3 Day Forecast
Climax (Nguyen et al., 2023) - 5.4 M 90.0 ± 5.69 58.1 ± 5.08 56.3 ± 5.12 66.8 ± 4.62 57.5 ± 5.30 83.4 ± 8.04 68.7 ± 2.35
FourCastNet (Pathak et al., 2022) FNO 5.3 M 88.7 ± 6.59 56.3 ± 4.83 54.4 ± 4.72 66.0 ± 4.40 55.2 ± 5.40 81.8 ± 8.49 67.1 ± 2.41
SFNONet (Bonev et al., 2023) SFNO 5.3 M 90.9 ± 5.32 58.0 ± 5.39 55.2 ± 5.28 67.2 ± 4.59 57.8 ± 5.66 83.4 ± 8.30 68.8 ± 2.40
SHNet (Ours) GSNO 4.0 M 92.5 ± 4.63 61.7 ± 4.68 60.8 ± 4.72 70.2 ± 4.21 63.6 ± 5.01 85.8 ± 7.46 72.4 ± 2.13
SFNONet (Bonev et al., 2023) GSNO 5.5 M 91.4 ± 4.97 59.8 ± 5.03 58.7 ± 4.95 68.4 ± 4.42 60.1 ± 5.40 84.3 ± 7.92 70.5 ± 2.27
SHNet SFNO 3.7 M 91.8 ± 4.82 60.6 ± 4.90 57.2 ± 5.06 68.7 ± 4.44 61.0 ± 5.22 84.6 ± 7.85 70.7 ± 2.24
@@ 5 Day Forecast
Climax (Nguyen et al., 2023) - 5.4 M 84.0 ± 9.63 28.2 ± 9.35 21.3 ± 8.23 35.7 ± 10.0 14.2 ± 8.57 66.3 ± 18.4 41.6 ± 4.59
FourCastNet (Pathak et al., 2022) FNO 5.3 M 83.0 ± 11.3 29.2 ± 8.68 22.5 ± 7.50 36.2 ± 10.5 13.8 ± 6.47 64.5 ± 19.9 41.5 ± 4.74
SFNONet (Bonev et al., 2023) SFNO 5.3 M 86.9 ± 7.76 35.8 ± 8.74 29.5 ± 9.28 45.3 ± 9.33 26.5 ± 9.47 71.2 ± 16.0 49.2 ± 4.27
SHNet (Ours) GSNO 4.0 M 88.2 ± 7.21 40.3 ± 8.32 35.2 ± 8.37 47.7 ± 8.71 30.7 ± 9.29 73.0 ± 14.8 52.5 ± 3.99
SFNONet (Bonev et al., 2023) GSNO 5.5 M 87.7 ± 7.68 38.6 ± 8.46 33.5 ± 8.72 47.1 ± 8.92 28.9 ± 9.35 72.1 ± 15.3 51.3 ± 4.11
SHNet SFNO 3.7 M 87.9 ± 7.40 38.2 ± 8.62 32.3 ± 8.85 46.3 ± 9.11 28.5 ± 9.36 72.4 ± 15.1 50.9 ± 4.10
5.2 Weather Forecasting
To further assess the performance of our methods, we utilize WeatherBench (Rasp et al., 2020), a widely used autoregressive global weather prediction benchmark that offers data at multiple spatial resolutions. We select the dataset with a spatial resolution of $5.625^{\circ}$ ( 32 × 64) and a temporal resolution of 1 hour for evaluation. The training period spans 1979-2015, validation is conducted on data from 2016, and testing covers 2017-2018. The full dataset includes 24 meteorological variables, and we select six key variables as prediction targets (Liu et al., 2024) (details in Appendix C.3). All models are under the same settings: batch size 128, 50 epochs, loss function of MRE, Adam optimizer with learning rate of $4\times 10^{-4}$. To evaluate multi-scale forecasting ability, we perform prediction at 24, 72, 120 autoregressive steps. Performance is measured using anomaly correlation coefficient (ACC) and latitude-weighted mean square error (MSE) (Nguyen et al., 2023) (details in Appendix C.3), reported in Table 2 and 8 (details in Appendix C.3). Besides, we conduct additional experiments at different resolution (64 × 128) and the results are summarized in the Table 6 in Appendix. Our method achieves the best results on all variables and time scales, especially in the predictions for the third and fifth days. The wind velocity predictions (V600) and their residual plots (darker plots are better) also demonstrate the accuracy and stability of our method in Figure 4.
To further evaluate superiority of the spectral kernel $G_{\theta_{2}}^{2}(l,m)$ as the spectral embedding in GSNO, we construct extra operator-level comparison experiments, with all other architectures and settings identical for fair comparison. The results in the Table LABEL:PE_vs clearly demonstrate that our GSNO consistently and significantly outperforms the SFNOs with higher capacity and spatial position-embedding SNO ($G_{\theta}(x,y)$) across all forecast lead times (details in Appendix C.3).
5.3 Diffusion MRI modelling of brain microstructure
Modeling brain microstructure with diffusion MRI is challenging due to sparse, anisotropic measurements across spherical shells (Van Essen et al., 2013; Jeurissen et al., 2014; Zeng et al., 2022); We choose dMRI-based Fiber Orientation Distribution (FOD) angular super-resolution (Zeng et al., 2022; Snoussi and Karimi, 2025) as the specific task for evaluating model performance on anisotropic system; see Appendix C.4 for challenge, motivation and preprocessing details. To further verify GSNO, we adopt the same architecture as FODNet (Zeng et al., 2022), only changing the convolutional layers, adapting it to SFNOs to obtain FOD-SFNO, and GSNOs to obtain FOD-GSNO. ACC is used for evaluation, which is Angular Correlation Coefficient, a standard metric for measuring the angular similarity between two spherical functions (Zeng et al., 2022). The results in Table 4 show that GSNO consistently outperforms other models, demonstrating strong generalization under irregular sampling and data sparsity. Figure 5 shows the region of interest taken from the cingulate gyrus. Compared with the obviously false positive predictions by other methods, the proposed GSNO achieves better prediction of fiber orientation and density.
Table 3: Operator-level comparison on WeatherBench. Bold is the best result.
@@ 5.3 Diffusion MRI modelling of brain microstructure
Methods SSMT-CSD (Khan et al., 2020) FOD-Net (Zeng et al., 2022) FOD-SFNO (Bonev et al., 2023) ESCNN (Snoussi and Karimi, 2025) FOD-GSNO (Ours)
@@ 5.4 Ablation Experiments and computational costs
To further evaluate the contributions of GSNO and SHNet, we conduct ablation studies to replace the GSNO block with a classic SFNO block in the SHNet or replace the classic SFNO block with GSNO block in the SFNONet, leading to performance degradation across all variables and time horizons (Table 2 and 8 in Appendix C.3). This confirms the effectiveness of GSNO and the multi-scale architecture of SHNet. Further, we observe more pronounced performance gains from GSNO blocks in longer-term predictions, suggesting that GSNO plays an increasingly important role in modelling long-range dynamics. To demonstrate the efficiency comparison, we provide the parameters of different models in Table 2. Detailed computational costs of different models are in Appendix C.6.
@@ 6 Conclusion
We have introduced a rigorous theoretical framework for spherical neural operators, grounded in the formulation of designable Green’s function on the sphere, and extended to the spherical harmonic domain. This framework is used to design different spherical operators and enables the principled incorporation of complex system constraint on spherical operator learning. Building on this framework, we design an absolute and relative position-dependent Green’s function, yielding a novel spherical operator solution beyond existing equivariant operator solution to capture complex physical constraints for asymmetric conditions and position-dependent phenomena. The proposed Green’s function leads to GSNO, designed to handle a broader range of real-world scenarios by the flexible balance between equivariance and invariance while retaining spherical geometry and grid invariance. Further, we design a multi-scale architecture, SHNet, incorporating geometrically adaptive up-and-down sampling to enhance interactions across resolutions on the sphere. Both theoretical analysis and extensive experiments on multiple spherical scenarios consistently demonstrate the superiority of our approaches over other state-of-the-art methods. This work lays the foundation for designing system-based operator solutions based on designable Green’s functions, suggesting a new pathway for the development of spherical modeling.